# Adaptive Gradient Normalization and Independent Sampling for (Stochastic) Generalized-Smooth Optimization

### v2

@@ 6.3.1 Proof outline and novelty
Furthermore, by leveraging the clipping structure of $h_{t}^{\beta}$ and Assumption 4, the left hand side can be lower bounded as $\frac{\gamma\|\nabla F(w_{t})\|^{2}}{h_{t}^{\beta}}\geq\min\{\gamma\|\nabla F(w_{t})\|^{2},\frac{\|\nabla F(w_{t})\|^{2-\beta}}{4((2\tau_{2}+2)/1-\tau_{1})^{\beta}L_{1}(2\tau_{1}^{2}+1)}\}.$ Finally, telescoping the above inequalities over $t$ and taking expectation leads to the desired bound in equation 17.
As a comparison, in the prior work on clipped SGD (Zhang et al., 2019; 2020a), their stochastic gradient and normalization term $h_{t}$ depend on the same mini-batch of samples, and therefore cannot be treated separately in the analysis. For example, their analysis proposed the following decomposition.
@@
Hence their analysis need to assume a constant upper bound for the approximation error $\|\nabla f_{\xi}(w_{t})-\nabla F(w_{t})\|$ in order to obtain a comparable bound to ours.
@@ 6.3.2 IAN-SGD under generalized PŁ condition
Under Assumption 2, we obtain the following recurion for IAN-SGD (See Appendix J for Proof details)
@@
Let Assumptions 1, 2, 3 and 4 hold. For the IAN-SGD algorithm, choose target error $0<\epsilon\leq\min\{1,1/2\mu\}$ and learning rate $\gamma=(2\mu\epsilon)^{\frac{4-2\beta}{\rho}}\cdot\min\{\frac{1}{4L_{0}(2\tau_{1}^{2}+1)},\frac{1}{4L_{1}((2\tau_{1}+1)/(1-\tau_{1}))(2\tau_{1}^{2}+1)},\frac{1}{8L_{1}(2\tau_{1}^{2}+1)(2\tau_{2}/(1-\tau_{1}))^{\beta}},\frac{L_{1}(2\tau_{1}^{2}+1)}{16((L_{0}+L_{1})(1+4\tau_{2}^{2})+L_{1}(\tau_{1}^{2}+1/2))^{2}}\}$, and $A=\frac{1}{1-\tau_{1}}$ $\delta=\frac{\tau_{2}}{1-\tau_{1}}$, $\Gamma=(4L_{1}\gamma(2\tau_{1}^{2}+1))^{\frac{1}{\beta}}$. Then, we have the following descent inequality
@@
We notice that equation 20 is almost the same as the recursion of AN-GD obtained in Appendix D. The main difference between equation 20 and equation 35 is the usage of $\gamma^{\frac{3}{2}}$ instead of $\gamma$ to characterize the recursion. As a result, the convergence rate of IAN-SGD is expected to be much slower than AN-GD, i.e., IAN-SGD converge with rate $\mathcal{O}((\frac{1}{\epsilon})^{\frac{3\beta}{\rho}}\log(\frac{1}{\epsilon}))$ when $\rho+\beta=2$; IAN-SGD converge with rate $\mathcal{O}((\frac{1}{\epsilon})^{\frac{3(2-\rho)}{\rho}})$ when $\rho+\beta<2$. We omit the discussions of proof as its convergence analysis exactly follows the proof presented in Appendix E.
@@ 7 Experiments
We conduct numerical experiments to compare IAN-SGD with other state-of-the-art stochastic algorithms, including the standard SGD (Ghadimi & Lan, 2013), normalized SGD (NSGD), clipped SGD (Zhang et al., 2019)(Clip SGD), SPIDER (Fang et al., 2018), normalized SGD with momentum (Cutkosky & Mehta, 2020)(NSGDm) etc. Compared with prior works that study clipped SGD (Zhang et al., 2019; 2020a; 2020b; Koloskova et al., 2023b; Reisizadeh et al., 2023), the main change of IAN-SGD lies in its usage of independent sampling and adaptive normalization. Thus, in experiment section, we evaluate convergence of algorithms in terms of total sample complexity. Additionally, to make empirical verification on adaptive normalization accordance with Theorem 1, our experiments focus on nonconvex phase retrieval, distributionally robust optimization (DRO), and deep neural network training. Specifically, prior works (Zhou et al., 2016; Liu et al., 2022a; Scaman et al., 2022) have shown that both phase retrieval and deep neural networks satisfy generalized PŁ conditions under mild assumptions. To align the DRO with our theoretical assumptions, we focus on regression tasks and adopt the $\chi^{2}$-divergence to model distributional uncertainty.
@@ 7.1 Nonconvex Phase Retrieval
The phase retrieval problem arises in optics, signal processing, and quantum mechanics (Drenth, 2007). The goal is to recover a signal from measurements where only the intensity is known, and the phase information is missing or difficult to measure. Specially, denote the underlying object as $x\in\mathbf{R}^{d}$. Suppose we take $m$ intensity measurements $y_{r}=|a_{r}^{T}x|^{2}+n_{r}$ for $r=1\cdots m$, where $a_{r}$ denotes the measurement vector and $n_{r}$ is the additive noise. We aim to reconstruct $x$ by solving the following regression problem.
@@
Such nonconvex function satisfies generalized-smooth with parameter $\alpha=\frac{2}{3}$. In this experiment, we generate the initialization $z_{0}\sim\mathcal{N}(1,6)$ and the underlying signal $x\sim\mathcal{N}(0,0.5)$ with dimension $d=100$. We take $m=3$k measurements with $a_{r}\sim\mathcal{N}(0,0.5)$ and $n_{r}\sim\mathcal{N}(0,4^{2})$. We implement all the stochastic algorithms in original form described in previous literatures. We set batch size $|B|=64$, and for IAN-SGD, we choose a small independent batch size $|B^{\prime}|=4$. We use fine-tuned learning rate for all algorithms, i.e., $\gamma=5e^{-5}$ for SGD, $\gamma=0.2$ for NSGD and NSGD with momentum, $\gamma=0.6$ for clipped SGD, and $\gamma=0.25$ for SPIDER and IAN-SGD. We set the maximal gradient clipping constant as $20$, $\delta=1e^{-3}$ for both clipped SGD and IAN-SGD. And we set normalization parameter $\beta=\frac{2}{3}$. Figure 2 (left) shows the comparison of objective value versus sample complexity. It can be observed that IAN-SGD consistently converges faster than other baseline algorithms.
@@ 7.2 Distributionally-Robust Optimization
Distributionally-robust optimization (DRO) is a popular approach to enhance robustness against data distribution shift. We consider the regularized DRO problem $\min_{w\in\mathcal{W}}f(w)=\sup_{\mathbb{Q}}\big{\{}\mathbb{E}_{\xi\sim\mathbb{Q}}[\ell_{\xi}(w)]-\lambda\Psi(\mathbb{P};\mathbb{Q})\big{\}},$ where $\mathbb{Q}$, $\mathbb{P}$ represents the underlying distribution and the nominal distribution respectively. $\lambda$ denotes a regularization hyper-parameter and $\Psi$ denotes a divergence metric. Under mild technical assumptions, Jin et al. (2021) showed that such a problem has the following equivalent dual formulation
$\min_{w\in\mathcal{W}}L(w,\eta)=\lambda\mathbb{E}_{\xi\sim P}\Psi^{*}\Big{(}\frac{\ell_{\xi}(w)-\eta}{\lambda}\Big{)}+\eta,$ (22)
where $\Psi^{*}$ denotes the conjugate function of $\Psi$ and $\eta$ is a dual variable. In particular, such dual objective function is generalized-smooth with parameter $\alpha=1$ (Jin et al., 2021; Chen et al., 2023). In this experiment, we use the life expectancy data (Arshi, 2017).
We set $\lambda=0.01$ and select $\Psi^{*}(t)=\frac{1}{4}(t+2)_{+}^{2}-1$, i.e., the conjugate of $\chi^{2}$-divergence. We adopt the regularized loss $\ell_{\xi}(\mathbf{w})=\frac{1}{2}(y_{\xi}-\mathbf{x}_{\xi}^{\top}\mathbf{w})^{2}+0.1\sum_{j=1}^{34}\ln(1+|w^{(j)}|)$.
For moving average parameter used for acceleration method, we set it as 0.1 and 0.25 for NSGD with momentum and SPIDER respectively. For stochastic algorithms without usage of multiple mini-batches, i.e., SGD, NSGD, NSGD with momentum and clipped SGD, we set their batch sizes as $|B|=128$. For SPIDER, we set $|B|=128$ and $|B^{\prime}|=2313$, where the algorithm will conduct a full-gradient computation after every 15 iterations. For IAN-SGD, we set the batch size for two batch samples as $|B|=128$ and $|B^{\prime}|=8$. We used fine-tuned learning rate for all algorithms, i.e., $\gamma=4e^{-5}$ for SGD, $\gamma=5e^{-3}$ for NSGD, NSGD with momentum and SPIDER, $\gamma=0.11$ for clipped SGD and IAN-SGD. We set the $\delta=1e^{-1}$, maximal gradient clipping constant as $30$, $25$ for clipped SGD and IAN-SGD respectively. And we set normalization parameter $\beta=\frac{2}{3}$. Figure 2 (right) shows the comparison of objective value versus sample complexity. It can be observed that objective value optimized by IAN-SGD consistently converges faster than other baselines algorithms.
@@ 7.3 Deep Neural Networks
According to Zhang et al. (2019), generalized-smooth has been observed to hold in deep neural networks. To further demonstrate the effectiveness of IAN-SGD algorithm, we conduct experiments for training deep neural networks. Specially, we train ResNet18, ResNet50 (He et al., 2016) on CIFAR10 Dataset (Krizhevsky, 2009) from scratch. We resize images as $32\times 32$ and normalize images with standard derivation equals to $0.5$ on each dimension. At the beginning of each algorithm, we fix the random seed and initialize model parameters using the Kaiming initialization. We compare our algorithm with several baseline methods, including SGD (Robbins & Monro, 1951), Adam (Kingma, 2014), Adagrad (Duchi et al., 2011a), NSGD, NSGD with momentum (Cutkosky & Mehta, 2020) and clipped SGD (Zhang et al., 2019).
For SGD, Adam and Adagrad, we utilize pytorch built-in optimizer to implement training pipeline. We implement training pipeline for NSGD, NSGD with momentum, clipped SGD and IAN-SGD. The normalization constant is computed through all model parameters at each iteration. The detailed algorithm settings are as following. For batch size, all algorithms use $B=128$, and $B^{\prime}=32$ for IAN-SGD. For moving average parameter, we use 0.9, 0.99 for Adam, and 0.25 for normalized SGD with momentum. For clipping threshold used in clipped SGD and IAN-SGD, we set them as 2 and $\delta=1e^{-1}$. The normalization power used for IAN-SGD is $\beta=\frac{2}{3}$. We use fine-tuned learning rate for all algorithms, i.e., $\gamma=1e^{-3}$ for SGD, Adam and Adagrad, $\gamma=1e^{-1}$ for NSGD and NSGD with momentum, $\gamma=2e^{-1}$ for clipped SGD and IAN-SGD. We trained ResNet18, ResNet50 on CIFAR10 dataset for 30 epochs and plot the training loss in Figure 3. Figure 3 (left) shows the training loss of ResNet18, Figure3 (right) shows the training loss of ResNet50. As we can see from these figures, the (pink) loss curve optimized by IAN-SGD indicates fast convergence rate comparable with several baselines, including SGD, NSGD NSGDm, clipped SGD, which demonstrate the effectiveness of IAN-SGD framework.
Based on experiments on non-convex phase retrieval, distributionally robust optimization (DRO), deep neural networks, and ablation study A.1, we found that IAN-SGD performs best when the adaptive normalization parameter is set to $\beta=\frac{2}{3}$. When $\beta$ is too small, i.e., $\beta\leq\frac{1}{2}$, IAN-SGD struggles to converge; when $\beta$ is close to 1, its performance resembles that of clipped SGD. To achieve fast convergence, the learning rate for IAN-SGD should lie between $1e^{-1}$ and $2e^{-1}$. Users should carefully fine-tune this parameter, as overly large learning rates may lead to divergence.

### v3

@@ 6.3.1 Proof outline and novelty
Furthermore, by leveraging the clipping structure of $h_{t}^{\beta}$ and Assumption 4, the left hand side of above inequality can be lower bounded as $\frac{\gamma\|\nabla F(w_{t})\|^{2}}{h_{t}^{\beta}}\geq\min\{\gamma\|\nabla F(w_{t})\|^{2},\frac{\|\nabla F(w_{t})\|^{2-\beta}}{4((2\tau_{2}+2)/1-\tau_{1})^{\beta}L_{1}(2\tau_{1}^{2}+1)}\}.$ Finally, telescoping the above inequalities over $t$ and taking expectation leads to the desired bound in equation 17.
As a comparison, in the previous works on clipped SGD (Zhang et al., 2019; 2020a), their stochastic gradient and normalization term $h_{t}$ depend on the same mini-batch of samples, and therefore cannot be treated separately in the analysis. For example, their analysis proposed the following decomposition.
@@
Hence their analysis needs to assume a constant upper bound for the approximation error $\|\nabla f_{\xi}(w_{t})-\nabla F(w_{t})\|$ in order to obtain a comparable bound to ours.
@@ 6.3.2 IAN-SGD under generalized PŁ condition
Under Assumption 2, we obtain the following recursion for IAN-SGD. See Appendix J for Proof details.
@@
Let Assumptions 1, 2, 3 and 4 hold. For the IAN-SGD algorithm, choose target error $0<\epsilon\leq\min\{1,1/2\mu\}$ and learning rate $\gamma=(2\mu\epsilon)^{\frac{4-2\beta}{\rho}}\cdot\min\{\frac{1}{4L_{0}(2\tau_{1}^{2}+1)},\frac{1}{4L_{1}((2\tau_{1}+2)/(1-\tau_{1}))(2\tau_{1}^{2}+1)},\frac{1}{8L_{1}(2\tau_{1}^{2}+1)(2\tau_{2}/(1-\tau_{1}))^{\beta}},\frac{L_{1}(2\tau_{1}^{2}+1)}{16((L_{0}+L_{1})(1+4\tau_{2}^{2})+L_{1}(\tau_{1}^{2}+1/2))^{2}}\}$, and $A=\frac{1}{1-\tau_{1}}$ $\delta=\frac{\tau_{2}}{1-\tau_{1}}$, $\Gamma=(4L_{1}\gamma(2\tau_{1}^{2}+1))^{\frac{1}{\beta}}$. Then, we have the following descent inequality
@@
We notice that equation 20 is almost the same as the recursion of AN-GD obtained in Appendix D. The main difference between equation 20 and equation 35 is the usage of $\gamma^{\frac{3}{2}}$ instead of $\gamma$ to characterize the recursion. As a result, the convergence rate of IAN-SGD is expected to be much slower than AN-GD, i.e., IAN-SGD converge with rate $\tilde{\mathcal{O}}({\epsilon}^{-\frac{3\beta}{\rho}})$ when $\rho+\beta=2$; IAN-SGD converge with rate $\mathcal{O}({\epsilon}^{-\frac{3(2-\rho)}{\rho}})$ when $\rho+\beta<2$. We omit the discussions of proof as its convergence analysis exactly follows the proof presented in Appendix E.
@@ 7 Experiments
We conduct numerical experiments to compare IAN-SGD with other state-of-the-art stochastic algorithms, including the standard SGD (Ghadimi & Lan, 2013), normalized SGD (NSGD), clipped SGD (Zhang et al., 2019), SPIDER (Fang et al., 2018), normalized SGD with momentum (Cutkosky & Mehta, 2020)(NSGDm) etc. Compared with previous works that study clipped SGD (Zhang et al., 2019; 2020a; 2020b; Koloskova et al., 2023; Reisizadeh et al., 2023), the main change of IAN-SGD lies in its usage of independent sampling and adaptive normalization. To empirically validate the effect of adaptive normalization in accordance with Theorem 1, our experiments focus on nonconvex phase retrieval, distributionally robust optimization (DRO), and deep neural network training. Prior works (Zhou et al., 2016; Jin et al., 2021; Liu et al., 2022a; Scaman et al., 2022) have shown that phase retrieval and deep neural networks satisfy generalized PŁ conditions under mild assumptions. Accordingly, we evaluate the convergence of all algorithms in terms of total sample complexity (for phase retrieval and DRO) and number of epochs (for deep neural networks). Furthermore, to align the DRO formulation with our theoretical assumptions, we focus on regression tasks and adopt the $\chi^{2}$-divergence to model distributional uncertainty.
@@ 7.1 Nonconvex Phase Retrieval
The phase retrieval problem arises in optics, signal processing, and quantum mechanics (Drenth, 2007). The goal is to recover a signal from measurements where only the intensity is known, and the phase information is missing or difficult to measure. Specially, denote the underlying object as $w\in\mathbf{R}^{d}$. Suppose we take $m$ intensity measurements $y_{r}=|a_{r}^{T}w|^{2}+n_{r}$ for $r=1\cdots m$, where $a_{r}$ denotes the measurement vector and $n_{r}$ is the additive noise. We aim to reconstruct $w$ by solving the following regression problem,
@@
Such nonconvex function satisfies generalized-smooth with parameter $\alpha=\frac{2}{3}$. In this experiment, we generate the initialization $w_{0}\sim\mathcal{N}(1,6)$ and the underlying true object $w^{*}\sim\mathcal{N}(0,0.5)$ with dimension $d=100$. We take $m=3$k measurements with $a_{r}\sim\mathcal{N}(0,0.5)$ and $n_{r}\sim\mathcal{N}(0,16)$. We implement all of the stochastic algorithms in original form as described in previous literature. We set batch size $|B|=64$, and for IAN-SGD, we choose a small independent batch size $|B^{\prime}|=4$. We fine-tuned learning rate for all algorithms, i.e., $\gamma=5e^{-5}$ for SGD, $\gamma=0.2$ for NSGD and NSGD with momentum, $\gamma=0.6$ for clipped SGD, and $\gamma=0.25$ for SPIDER and IAN-SGD. We set the maximal gradient clipping constant as $20$, $\delta=1e^{-3}$ for both clipped SGD and IAN-SGD. And we set normalization parameter $\beta=\frac{2}{3}$. Figure 2 (left) shows the comparison of objective value versus sample complexity. It can be observed that IAN-SGD consistently converges faster than other baseline algorithms.
@@ 7.2 Distributionally-Robust Optimization
Distributionally-robust optimization (DRO) is a popular approach to enhance robustness against data distribution shift. We consider the regularized DRO problem $\min_{w\in\mathbf{R}^{d}}f(w)=\sup_{\mathbb{Q}}\big{\{}\mathbb{E}_{\xi\sim\mathbb{Q}}[\ell_{\xi}(w)]-\lambda\phi(\mathbb{P};\mathbb{Q})\big{\}},$ where $\mathbb{Q}$, $\mathbb{P}$ represents the underlying distribution and the nominal distribution respectively. $\lambda$ denotes a regularization hyper-parameter and $\phi$ denotes $\phi$-divergence metric. Under mild technical assumptions, Jin et al. (2021) showed that such a problem has the following equivalent dual formulation
$\min_{w\in\mathbf{R}^{d},\eta\in\mathbf{R}}L(w,\eta)=\lambda\mathbb{E}_{\xi\sim P}\phi^{*}\Big{(}\frac{\ell_{\xi}(w)-\eta}{\lambda}\Big{)}+\eta,$ (22)
where $\phi^{*}$ denotes the convex conjugate function of $\phi$. In particular, such dual objective function satisfies the generalized-smooth condition 2 with parameter $\alpha=1$ (Jin et al., 2021; Chen et al., 2023). In this experiment, we use the Life Expectancy data (Arshi, 2017) for regression task. For objective equation 22, we set $\lambda=0.01$, select $\phi^{*}(t)=\frac{1}{4}(t+2)_{+}^{2}-1$, which is the convex conjugate of $\chi^{2}$-divergence. We adopt the regularized loss $\ell_{\xi}({w})=\frac{1}{2}(y_{\xi}-{x}_{\xi}^{\top}{w})^{2}+0.1\sum_{j=1}^{34}\ln(1+|w^{(j)}|)$, where $x_{\xi},y_{\xi}$ represent input and output data, respectively. And we optimize the primal and dual variable of $L(w,\eta)$ in parallel following conclusion drawn from Jin et al. (2021).
For the stochastic momentum moving average parameter used for acceleration methods, we set it as $0.1$ and $0.25$ for NSGD with momentum and SPIDER respectively. For stochastic algorithms without usage of multiple mini-batches, i.e., SGD, NSGD, NSGD with momentum and clipped SGD, we set their batch sizes as $|B|=128$. For SPIDER, we set $|B|=128$ and $|B^{\prime}|=2313$, where the algorithm will conduct a full-gradient computation after every $15$ iterations. For IAN-SGD, we set the batch size for two batch samples as $|B|=128$ and $|B^{\prime}|=8$. We fine-tuned the learning rate for each algorithm, i.e., $\gamma=4e^{-5}$ for SGD, $\gamma=5e^{-3}$ for NSGD, NSGD with momentum and SPIDER, $\gamma=0.11$ for clipped SGD and IAN-SGD. We set the $\delta=1e^{-1}$, maximal gradient clipping constant as $30$, $25$ for clipped SGD and IAN-SGD respectively. And we set normalization parameter $\beta=\frac{2}{3}$. Figure 2 (right) shows the comparison of objective value versus sample complexity. It can be observed that objective value optimized by IAN-SGD consistently converges faster than other baselines algorithms.
@@ 7.3 Deep Neural Networks
According to Zhang et al. (2019), generalized-smooth condition 2 has been observed to hold in deep neural networks. To further demonstrate the effectiveness of IAN-SGD algorithm, we conduct experiments for training deep neural networks. Specially, we train ResNet18 and ResNet50 (He et al., 2016) on CIFAR10 Dataset (Krizhevsky, 2009) from scratch. We resize images as $32\times 32$ and normalize images with standard derivation equals to $0.5$ on each dimension. At the beginning of each algorithm, we fix the random seed and initialize model parameters using the Kaiming initialization (He et al., 2015). We compare our algorithm with several baseline methods, including SGD (Robbins & Monro, 1951), Adam (Kingma, 2014), Adagrad (Duchi et al., 2011a), NSGD, NSGD with momentum (Cutkosky & Mehta, 2020) and clipped SGD (Zhang et al., 2019).
For SGD, Adam and Adagrad, we utilize PyTorch built-in optimizer (Paszke et al., 2019) to implement training pipelines. We implement training pipelines for NSGD, NSGD with momentum, clipped SGD and IAN-SGD. The normalization constant is computed through all model parameters at each iteration. The detailed algorithm settings are as follows. For batch size, all algorithms use $B=128$, and $B^{\prime}=32$ for IAN-SGD. For stochastic momentum moving average parameter, we set it as $0.9$, $0.99$ for Adam, and 0.25 for normalized SGD with momentum. For clipping threshold used in clipped SGD and IAN-SGD, we set them as 2 and $\delta=1e^{-1}$. The normalization power used for IAN-SGD is $\beta=\frac{2}{3}$. We fine-tuned learning rate for all algorithms, i.e., $\gamma=1e^{-3}$ for SGD, Adam and Adagrad, $\gamma=1e^{-1}$ for NSGD and NSGD with momentum, $\gamma=2e^{-1}$ for clipped SGD and IAN-SGD. We trained ResNet18, ResNet50 on CIFAR10 dataset for 30 epochs and plot the training loss in Figure 3. Figure 3 (left) shows the training loss of ResNet18, Figure3 (right) shows the training loss of ResNet50. As we can see from these figures, the (pink) loss curve optimized by IAN-SGD indicates fast convergence rate comparable with several baselines, including SGD, NSGD NSGDm, clipped SGD, which demonstrate the effectiveness of IAN-SGD framework.
Based on our experiments on non-convex phase retrieval, distributionally robust optimization (DRO), deep neural networks, and an ablation study A.1, we found that IAN-SGD performs best when the adaptive normalization parameter is set to $\beta=\frac{2}{3}$. When $\beta$ is too small, i.e., $\beta\leq\frac{1}{2}$, IAN-SGD struggles to converge; when $\beta$ is close to 1, its performance resembles that of clipped SGD. To achieve fast convergence, the learning rate for IAN-SGD should lie between $1e^{-1}$ and $2e^{-1}$. Users should carefully fine-tune this parameter, as overly large learning rates may lead to divergence.
@@ 8 Conclusion
9 Acknowledgements
We are grateful to the anonymous reviewers and Dr. Ziyi Chen for insightful comments that led to several improvements in the presentation and technical details. This research was supported in part by the Air Force Research Laboratory Information Directorate (AFRL/RI), through AFRL/RI Information Institute®, contract number FA8750-20-3-1003. The work of Yufeng Yang and Yi Zhou was supported by the National Science Foundation under grants DMS-2134223, ECCS-2237830. The work of Shaofeng Zou was supported in part by National Science Foundation under Grants CCF-2438429 and ECCS-2438392(CAREER). The work of Yifan Sun was supported in part by the ONR grant W911NF-22-1-0292.